# CAGE: Curvature-Aware Gradient Estimation For Accurate Quantization-Aware Training

### v1

@@ 5 Conclusion and Discussion
The key theoretical contribution is the definition of Pareto-optimal solutions for quantized optimization. Moreover, CAGE offers strong ergodic convergence guarantees to a Pareto-optimal point in the smooth non-convex setting, providing theoretical grounding often lacking in STE heuristics. While the approach is optimizer-agnostic, we developed a highly efficient, decoupled implementation on top of AdamW that leverages existing optimizer statistics for stability and performance. Our formulation also generalizes the concurrent work of [18].
@@
The success of CAGE indicates that accurately estimating the gradient by incorporating local curvature information is crucial for advancing QAT performance. By explicitly counteracting the impact of quantization error on the optimization dynamics, CAGE provides a robust and theoretically sound framework for training accurate low-bit models, helping to bridge the gap between quantized efficiency and full-precision performance. Future work may explore the application of CAGE in ultra-low-bit regimes and its integration with other compression techniques, such as sparsification and vector quantization.
@@ Appendix B Hyperparameters and Reproducibility
Shared settings. Sequence length L=512L{=}512; batch size $64$ with gradient accumulation $8$ (effective tokens/step fixed across sizes). Optimizer AdamW with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$, $\varepsilon{=}10^{-8}$, weight decay $0.1$, gradient clip $1.0$. Cosine LR schedule with warmup $10\%$ of total steps. FP32 master weights and optimizer states; bfloat16 compute. Hardware: $8\times$ H100 (NCCL, compile enabled).
@@
CAGE settings. Decoupled variant. silence ratio s=0.9s{=}0.9. $\lambda{=}2.0$

### v2

@@ 5 Summary and Discussion
The key theoretical contribution is the definition of Pareto-optimal solutions for quantized optimization. Moreover, CAGE offers strong ergodic convergence guarantees to a Pareto-optimal point in the smooth non-convex setting, providing theoretical grounding often lacking in STE heuristics. While the approach is optimizer-agnostic, we developed a highly efficient, decoupled implementation on top of AdamW that leverages existing optimizer statistics for stability and performance. Our formulation also generalizes the concurrent work of kwun2025lotion.
@@
The good performance of CAGE indicates that incorporating local curvature information and regularization are useful for QAT performance. Future work may explore the application of CAGE in ultra-low-bit regimes and its integration with other compression techniques, such as sparsification and vector quantization.
@@ Appendix B Hyperparameters and Reproducibility
Shared settings. Sequence length $L=512$; batch size $64$ with gradient accumulation $8$ (effective tokens/step fixed across sizes). Optimizer AdamW with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\varepsilon=10^{-8}$, weight decay $0.1$, gradient clip $1.0$. Cosine LR schedule with warmup $10\%$ of total steps. FP32 master weights and optimizer states; bfloat16 compute. Hardware: $8\times$ H100 (NCCL, compile enabled).
@@
CAGE settings. Throughout the experiments, unless otherwise stated, we use the decoupled variant, with silence ratio $s=0.9$, and regularization parameter $\lambda=2.0$.
Figure 9: Validation perplexity for CAGE under different choices of silence ratio $s$ and regularization coefficient $\lambda$ on the 50M Llama model with W4A4 quantization.
Appendix C Ablation of CAGE Hyperparameters
We conduct a two-dimensional sweep over the CAGE hyperparameters; the silence ratio $s$ and the regularization coefficient $\lambda$. Figure 9 reports validation perplexity for runs on a 50M Llama model trained on C4 in the W4A4 setting. We find that $s\in[0.8,0.95]$ and $\lambda\in[1,2.5]$ form a reliable operating region. Within this range, CAGE consistently improves over the baseline.